# V2X Cooperative Perception for Autonomous Driving: Recent Advances and Challenges

### v4

@@ IV-E4 Hybrid Fusion
Objects near the onboard sensor typically exhibit a high point density and are more easily detected using data from a single sensor, while distant points tend to be sparse due to limited sensor visibility. Early collaboration for distant regions can improve detection accuracy, as demonstrated by research [220], which proposed a V2I cooperative 3D object detection model using hybrid fusion. This method transmits raw data (early fusion) for low-density areas and shares object-level information (late fusion) for nearby areas, providing a cost-effective solution but with limited flexibility in adjusting data sharing. Hybrid CP [222] proposed a region-based hybrid collaborative perception strategy. Collaboration is classified into two types based on the overlap between the detection ranges of vehicles. For the overlapping area, feature-level intermediate collaboration is used by exchanging and fusing features from various CAVs. Late collaboration is carried out in the non-overlapping area by developing and sharing the local detection result with a low communication volume. Conversely, ML-Cooper [221] employs an improved Soft Actor-Critic [265] algorithm to optimize V2V bandwidth by dynamically adjusting data sharing among raw, feature, and object levels based on channel conditions, enhancing adaptability in dynamic vehicular networks.
@@ V Issues of Cooperative Perception and Corresponding Methods
While the majority of research on CP for driving automation has focused on key aspects of the collaboration process, such as communication latency, information alignment, and fusion strategies, as discussed in the previous section, limited attention has been given to subtle yet crucial challenges that have practical implications for advancing Cooperative ITS. This section delves into exploring recent findings and the corresponding methodologies to address these practical challenges. Table VI provides a comprehensive overview of characteristics related to these advancements.
@@ V-A1 Model Heterogeneity
Object-level CP frameworks avoid the need for full model data exchange by sharing perception results like bounding boxes and confidence scores participating agents [225]. Nevertheless, due to model heterogeneity, the confidence scores provided by different agents may not always align accurately. In [225], a model-agnostic CP framework was proposed to tackle this confidence misalignment problem, using a Doubly Bounded Scaling calibrator to eliminate bias in confidence scores and a Promote-suppress Aggregation algorithm to select high-score bounding boxes based on calibrated confidence and spatial consistency. In addition to variations in confidence levels, connected agents’ detection outcomes may vary due to differences in motion prediction models, with some agents lacking information like acceleration or yaw rate. To address heterogeneous perception fusion with varying state spaces, a CI-based track-to-track fusion approach was proposed in [226], augmenting low-dimensional tracks and exploring methods for state augmentation.
Late collaboration preserves model privacy by allowing agents to train models individually and share outputs, but this can introduce noise and degrade performance. Intermediate collaboration is more flexible and popular for V2X CP, but feature extractors’ differences can create a domain gap. To overcome this issue, a Multi-agent Perception Domain Adaptation (MPDA) framework was introduced in [227], incorporating a feature resizer and a sparse cross-domain transformer to align and produce domain-invariant features, thereby minimizing disparities between agents utilizing heterogeneous feature extraction networks. Inspired by the parameter-efficient fine-tuning strategy [266], MACP [228] introduced a Convolution Adapter (ConAda) to mitigate domain shifts by projecting the source feature map to a lower dimension and then transforming it back. Additionally, a single-agent pre-trained model was adapted to cooperative settings by freezing most parameters and adding lightweight modules. While this approach effectively mitigates domain shifts and enhances perception, it may exhibit reduced precision for objects in the 50-100m range due to inherent biases in the pre-trained model.
@@ V-A2 Data Heterogeneity
In a multi-agent hetero-modal setting, connected agents have various sensors, forming a dynamic heterogeneous graph with stochastic sensor types and changing poses across scenes. To handle these challenges, HM-ViT [164] introduced the H3GAT module for V2V cooperative 3D detection, which updates node states by extracting inter-agent and intra-agent cues using transformer blocks and a hetero-modal MLP layer. This allows processing camera and LiDAR features with distinct parameters, enabling flexibility, robustness, and superior performance in scenarios with diverse sensor modalities.
@@
On the other hand, identical modality sensors can vary in brand, resolution, and configuration, leading to domain discrepancies that impact system performance. In response to this exigency, DI-V2X [229] introduced a teacher-student distillation architecture with three key components: the Domain-Mixing Instance Augmentation module for aligning distributions between teacher and student models, the Domain-Adaptive Fusion module for combining vehicle and infrastructure features, and the Progressive Domain-Invariant Distillation module to align student and teacher features before and after fusion. This method improves V2X perception by creating a generalized representation across various domains.
@@
Besides sensor characteristics, diverse environmental conditions inevitably create a substantial domain gap and data heterogeneity among CAVs [230]. To address this, [230] proposed a domain generalization framework with an amplitude augmentation module to reduce the model’s influence on low-spectral variations, a domain alignment mechanism to unify image pixel distributions, and a meta-consistency training scheme to simulate domain shifts and optimize the model with consistency loss, potentially eliminating the domain discrepancy among CAVs prior to inference.
@@
The distinctness in data characteristics across different agent types presents significant challenges in CP. VINet [231] addressed this challenge using a Two-stream Fusion algorithm to integrate heterogeneous features from infrastructures and CAVs via a feature mapping and max-out procedure on an edge server. In V2X-M2C [232], a heterogeneity-reflected convolution module was introduced to manage multiple heterogeneous agents using two parallel convolution pipelines, considering their different feature distributions while maintaining a lightweight and efficient architecture. However, time asynchrony and calibration errors between vehicle and infrastructure cameras can lead to inaccurate relative position detection, causing misalignment between ground truth and projected 2D bounding boxes. VIMI [233] tackled the calibration noise issue through its Multi-scale Cross Attention and Camera-aware Channel Masking modules, using multi-scale feature fusion and camera parameter priors to correct for misalignment caused by time asynchrony.
@@ V-A3 Model and Data Heterogeneity
While some existing studies address the heterogeneity of data and models, effectively modelling dynamically increasing heterogeneous scenarios continues to pose significant challenges. In reality, new heterogeneous agent types with different modalities and models may continuously emerge, causing domain gaps and limiting scalability. HEAL [234] solved this problem by establishing a unified feature space for the initial homogeneous agents and then aligning new heterogeneous agents to this space through a novel reverse alignment mechanism. This approach enables the integration of new agent types with minimal training costs while preserving model and data privacy. Despite its scalability and efficiency, HEAL’s reliance on BEV features requires compatibility with keypoint-based LiDAR models, and its performance depends on accurate agent localization, which can be affected by localization noise in real-world scenarios.
@@ V-B Lossy Communication
In practical V2X communication, factors such as interference, tampering, Doppler shifts [267], routing failures [268], multipath effects, and signal fluctuations [269] frequently cause incomplete or inaccurate information exchange, thereby diminishing the performance of CP.
Drawing inspiration from image-denoising networks, the Lossy Communication Aware Repair Network (LCRN) [235] utilized an encoder-decoder architecture with skip connections to generate tensor-wise filters for repairing damaged features from lossy communication. These filters are applied to lossy features, which are then fused via a V2V attention module. While LCRN mitigates the impact of feature degradation, it does not directly address feature loss. In contrast, V2X-INCOP [236] recovered missing data by leveraging historical information, using an adaptive communication model that extracts multi-scale spatial-temporal features to estimate missing information, effectively countering communication interruptions. In [243], a Historical Frame Prediction (HFP) module was introduced to address information loss or delay issues by predicting the current frame with temporal associations, taking the stacked results of the previous few historical frames as input.
Previous approaches often overlook realistic channel models and may not capture the complexity of real-world communication channels. To address this, [237] incorporated Rician fading with free-space path loss and the WINNER II model [270] with multi-path fading to simulate practical channels. A self-supervised adaptive weighting model, based on a CNN with a Softmax layer, was introduced to mitigate channel distortion effects on intermediate fusion. The model was trained using contrastive self-supervision without manual annotations, and its generalization was validated across various data domains, detection backbones, noise levels, and path loss factors.
Accessing historical information can reduce transmission frequency and provide a compensatory mechanism for potential communication failures, thereby improving system reliability and performance. Building on this concept, TempCoBEV [238] introduced an importance-guided attention architecture to prioritize critical regions from the historical context, compensating for communication failures. This temporal module can be integrated into state-of-the-art camera-based CP models without requiring complete model re-training, significantly reducing training time. Experiments on the OPV2V dataset demonstrate that TempCoBEV outperforms non-temporal models in predicting current and future BEV map segmentations, particularly in scenarios involving communication failures.
@@ V-C Communication Delays
Latency resulting from network congestion reduces perception accuracy and disrupts fusion alignment [271]. A basic solution to address latency issues is to adjust the timestamp by compensating for the time difference with a reference agent. Building on this concept, CooperFuse [239] introduced a multi-agent temporal synchronization module, incorporating time calibration communication, time compensation, and calibration completion to minimize network transmission latency.
The time series prediction strategy is also effective in addressing the latency issue by predicting the sender’s perception at the current time step based on historical frames. In this context, the FFNet [240] leveraged temporal coherence by extracting the first-order derivative of feature flow from previous infrastructure frames to predict future features and mitigate asynchrony. While effective for very short time intervals, this first-order expansion method has limitations. SyncNet [241] adopted a dual-branch pyramid Long Short-term Memory (LSTM) network to infer real-time features and cooperative attention. The HFP module [243] can generate the prediction of the current frame through a series of multidimensional convolution mixing operations. However, these approaches have not effectively addressed issues caused by irregular time delays. As a substitute, CoBEVFlow [242] treated each collaboration message as an irregular sample, using historical frames to estimate the BEV flow map for asynchronous feature alignment, though it still depends on past data for latency compensation. In contrast, V2X-PC [136] bypassed this reliance by directly utilizing low-level coordinate information in point clusters to predict their locations at the current timestamp.
Re-encoding received features with delay time is another approach for mitigating communication delays. For instance, V2X-ViT [145] addressed time delay through a Delay-aware Positional Encoding module that utilizes sinusoid functions conditioned on delay time and channel information to initialize a learnable embedding. Similarly, V2VNet [140] utilized a CNN to manage latency by incorporating the received intermediate feature, relative 6DoF pose, and delay time, producing a time-delay-compensated representation. These methods assume latency is known; however, latency is typically calculated using timestamps from different agents’ clocks, which may result in deviations and inaccurate latency measurements. To address this, FreeAlign [195] introduced a multi-anchor based subgraph searching algorithm to identify the common subgraph between salient-object graphs across different agents, estimating clock deviations and determining the precise time difference between collaborative messages. This approach ensures accurate temporal alignment without relying on external clock signals.
@@ V-D1 Data Security
While cooperative multi-agent systems offer significant promise, they also present security risks due to the potential for malicious or unreliable communication between agents [272]. Malicious agents may tamper with shared packets or generate forged data. The ego vehicle is unable to independently verify missed events or confirm the authenticity of received data [273]. Imperceptible perturbations introduced by adversaries can substantially alter perception outputs, compromising system integrity. Although adversarial training has been used to address these threats [244], it adds training overhead and often fails to generalize to novel attacks [274]. As a solution, ROBOSAC [245] was proposed, allowing ego vehicles to intelligently select trustworthy collaborators. By detecting divergences in adversarial messages and aligning with reliable collaborators, ROBOSAC generalizes to unseen adversarial strategies and offers a flexible trade-off between performance and computational cost.
@@
To ensure data security, assessing the quality of received information and detecting erroneous data before processing is crucial. Probabilistic modelling techniques have been utilized to estimate the trustworthiness of connected agents based on data consistency [246]. This approach applies a Bayes filter to quantify consistency, allowing the detection of forged information by analyzing the correlation between data consistency and agent reliability. However, this method is constrained in detecting attackers, as it requires substantial redundancy—specifically, at least triple redundancy in conflicting data—to effectively identify malicious agents.
@@
FDII [247] addressed spoofing attacks by leveraging LiDAR data from neighbouring vehicles to detect tampered points in the victim’s scan. By comparing the altered and unaltered scans, FDII identifies discrepancies and classifies attacks using a decision tree algorithm, which also updates the unsafe areas for each vehicle. The approach proved effective in identifying attack types and reconstructing vulnerable regions. However, decision trees may struggle with complex data relationships and often a bias toward the majority class in imbalanced datasets, potentially leading to the suboptimal classification of minority classes.
@@
To counter data fabrication attacks, the Collaborative Anomaly Detection (CAD) model [248] allows benign vehicles to collectively detect malicious fabrications by sharing and merging occupancy maps. These maps, represented by fine-grained polygons, offer greater precision and flexibility compared to grid-based alternatives. CAD incorporates motion estimation to track objects across frames, attaching motion data to the maps. Consistency checks are performed to identify discrepancies among synchronized maps and compare perception results with the merged map, raising alerts for conflicted regions that suggest potential anomalies. Although CAD proves effective in real-world scenarios, its efficacy depends on the presence of at least one benign CAV monitoring the attacked region.
@@
Malicious Agent Detection (MADE) [249] used two key statistics, match loss and collaborative reconstruction loss, to identify malicious agents. Match loss quantifies discrepancies in bounding box proposals between the ego agent and the inspected agent, while collaborative reconstruction loss measures the consistency of feature maps when the two agents collaborate. These metrics can be applied independently or together to form a multi-test with a controlled false detection rate, utilizing the Benjamini-Hochberg procedure [275]. Although this approach offers valuable insights, MADE’s reliance on conformal p-values computed from calibration sets assumes the ego agent is benign, a condition that may not always be valid in real-world scenarios.

### v5

@@ IV-E4 Hybrid Fusion
Objects near the onboard sensor typically exhibit a high point density and are more easily detected using data from a single sensor, while distant points tend to be sparse due to limited sensor visibility. Early collaboration for distant regions can improve detection accuracy, as demonstrated by research [133], which proposed a V2I cooperative 3D object detection model using hybrid fusion. This method transmits raw data (early fusion) for low-density areas and shares object-level information (late fusion) for nearby areas, providing a cost-effective solution but with limited flexibility in adjusting data sharing. In Hybrid CP [135], a region-based hybrid collaborative perception strategy was proposed, in which collaboration was classified into two types based on the overlap between the detection ranges of vehicles. For overlapping areas, feature-level intermediate collaboration was performed by exchanging and fusing features from multiple CAVs. In contrast, late collaboration was employed in non-overlapping areas by generating and sharing local detection results with low communication volume. To enhance adaptability in dynamic vehicular networks, ML-Cooper [134] was developed to optimize V2V bandwidth by dynamically adjusting data sharing across raw, feature, and object levels according to current channel conditions.
@@ V Issues of Cooperative Perception and Corresponding Methods
While the majority of research on CP for driving automation has focused on key aspects of the collaboration process, such as communication latency, information alignment, and fusion strategies, as discussed in the previous section, limited attention has been given to subtle yet crucial challenges that have practical implications for advancing Cooperative ITS. This section delves into exploring recent findings and the corresponding methodologies to address these practical challenges. Table V provides a comprehensive overview of characteristics related to these advancements.
@@ V-A1 Model Heterogeneity
Object-level cooperative perception frameworks eliminate the need for full model data exchange by sharing perception outputs such as bounding boxes, confidence scores, and participating agents [138]. However, due to model heterogeneity, the confidence scores generated by different agents may not always align accurately. To address this issue, a model-agnostic cooperative perception framework was proposed in [138], in which a Doubly Bounded Scaling calibrator was used to eliminate bias in the confidence scores, and a Promote-suppress Aggregation algorithm was employed to select high-confidence bounding boxes based on calibrated scores and spatial consistency.
In addition to variations in confidence levels, differences in motion prediction models may lead to discrepancies in detection outcomes among connected agents, particularly when some agents lack information such as acceleration or yaw rate. To address heterogeneous perception fusion with varying state spaces, a CI-based track-to-track fusion approach was proposed in [139], augmenting low-dimensional tracks and exploring methods for state augmentation.
Late collaboration preserves model privacy by enabling agents to train models individually and share only their outputs. However, this approach may introduce noise and degrade performance. Intermediate collaboration is more flexible and commonly adopted in V2X CP, but differences in feature extractors can result in a domain gap. To address this issue, a Multi-agent Perception Domain Adaptation (MPDA) framework was introduced in [140], in which a feature resizer and a sparse cross-domain transformer were incorporated to align and produce domain-invariant features, thereby minimizing disparities between agents utilizing heterogeneous feature extraction networks. In [141], a Convolution Adapter (ConAda) was introduced to mitigate domain shifts by projecting the source feature map to a lower dimension and subsequently transforming it back. Furthermore, a single-agent pre-trained model was adapted to cooperative settings by freezing most parameters and integrating lightweight modules. Although this approach effectively mitigated domain shifts and enhanced perception, reduced precision for objects within the 50–100m range was observed, attributed to inherent biases in the pre-trained model.
@@ V-A2 Data Heterogeneity
In a multi-agent hetero-modal setting, connected agents have various sensors, forming a dynamic heterogeneous graph with stochastic sensor types and changing poses across scenes. To address these challenges, HM-ViT [75] was proposed, in which the H3GAT module was introduced for vehicle-to-vehicle cooperative three-dimensional object detection. In this framework, node states were updated by extracting inter-agent and intra-agent cues using transformer blocks and a hetero-modal MLP layer. This design allowed camera and LiDAR features to be processed with distinct parameters, enabling greater flexibility, robustness, and improved performance in scenarios involving diverse sensor modalities.
@@
On the other hand, identical modality sensors can vary in brand, resolution, and configuration, leading to domain discrepancies that impact system performance. In response to this challenge, DI-V2X [142] was proposed, in which a teacher-student distillation architecture was introduced, consisting of three key components: the Domain-Mixing Instance Augmentation module was designed to align data distributions between the teacher and student models, the Domain-Adaptive Fusion module was employed to combine vehicle and infrastructure features, and the Progressive Domain-Invariant Distillation module was used to align features from the student and teacher models both before and after fusion. Through this design, a generalized representation across various domains was achieved, thereby improving V2X perception performance.
@@
Besides sensor characteristics, diverse environmental conditions inevitably introduce a substantial domain gap and data heterogeneity among CAVs [143]. To address this, a domain generalization framework was proposed in [143], in which an amplitude augmentation module was employed to reduce the model’s sensitivity to low-spectral variations, a domain alignment mechanism was used to unify image pixel distributions, and a meta-consistency training scheme was adopted to simulate domain shifts and optimize the model using consistency loss. This approach has the potential to eliminate domain discrepancies among CAVs prior to inference.
@@
The distinctness in data characteristics across different agent types presents significant challenges in CP. This issue was addressed in VINet [144], where a Two-stream Fusion algorithm was employed to integrate heterogeneous features from infrastructures and CAVs through a feature mapping and max-out procedure performed on an edge server. In V2X-M2C [145], a heterogeneity-reflected convolution module was introduced to manage multiple heterogeneous agents using two parallel convolution pipelines, considering their different feature distributions while maintaining a lightweight and efficient architecture. However, time asynchrony and calibration errors between vehicle and infrastructure cameras can lead to inaccurate relative position detection, causing misalignment between ground truth and projected 2D bounding boxes. To address this, VIMI [146] was proposed, in which calibration noise was mitigated through the use of Multi-scale Cross Attention and Camera-aware Channel Masking modules. These modules applied multi-scale feature fusion and incorporated camera parameter priors to correct for misalignment caused by time asynchrony.
@@ V-A3 Model and Data Heterogeneity
While some existing studies address the heterogeneity of data and models, effectively modelling dynamically increasing heterogeneous scenarios continues to pose significant challenges. In reality, new heterogeneous agent types with different modalities and models may continuously emerge, causing domain gaps and limiting scalability. This issue was addressed in HEAL [147], where a unified feature space was first established for the initial homogeneous agents, and new heterogeneous agents were subsequently aligned to this space using a novel reverse alignment mechanism. Through this approach, the integration of new agent types was enabled with minimal training costs while preserving both model and data privacy. Despite its scalability and efficiency, the reliance of HEAL on BEV features necessitated compatibility with keypoint-based LiDAR models, and its performance was shown to be dependent on accurate agent localization, which may be affected by localization noise in real-world scenarios.
@@ V-B Lossy Communication
In practical V2X communication, factors such as interference, tampering, Doppler shifts, routing failures, multipath effects, and signal fluctuations frequently cause incomplete or inaccurate information exchange, thereby diminishing the performance of CP.
Drawing inspiration from image-denoising networks, the Lossy Communication Aware Repair Network (LCRN) [148] utilized an encoder-decoder architecture with skip connections to generate tensor-wise filters for repairing damaged features from lossy communication. These filters are applied to lossy features, which are then fused via a V2V attention module. While LCRN mitigates the impact of feature degradation, it does not directly address feature loss. In contrast, V2X-INCOP [149] was developed to recover missing data by leveraging historical information through an adaptive communication model that extracted multi-scale spatial-temporal features to estimate missing information, thereby countering communication interruptions. Additionally, in [156], a Historical Frame Prediction (HFP) module was introduced to address information loss or delay issues by predicting the current frame with temporal associations, using the stacked results of the previous few historical frames as input.
Previous approaches often overlooked realistic channel models and failed to capture the complexity of real-world communication environments. To address this limitation, Rician fading with free-space path loss and the WINNER II model [178], incorporating multi-path fading, were adopted in [150] to simulate practical communication channels. A self-supervised adaptive weighting model, based on a CNN with a Softmax layer, was introduced to mitigate channel distortion effects on intermediate fusion. The model was trained using contrastive self-supervision without manual annotations, and its generalization was validated across various data domains, detection backbones, noise levels, and path loss factors.
Access to historical information has been shown to reduce transmission frequency and provide a compensatory mechanism for potential communication failures, thereby enhancing system reliability and performance. Building on this concept, an importance-guided attention architecture was introduced in TempCoBEV [151] to prioritize critical regions from the historical context, compensating for communication failures. This temporal module was designed to be integrated into state-of-the-art camera-based CP models without requiring complete model re-training, resulting in a significant reduction in training time. Experimental results on the OPV2V dataset demonstrated that TempCoBEV outperformed non-temporal models in predicting current and future BEV map segmentations, particularly in scenarios involving communication failures.
@@ V-C Communication Delays
Latency caused by network congestion has been shown to reduce perception accuracy and disrupt fusion alignment [179]. A basic solution to mitigate latency involves adjusting timestamps by compensating for the time difference relative to a reference agent. Building on this concept, a multi-agent temporal synchronization module was introduced in CooperFuse [152], in which time calibration communication, time compensation, and calibration completion were incorporated to minimize the effects of network transmission latency.
The time series prediction strategy has been demonstrated to be effective in addressing latency by estimating a sender’s perception at the current time step based on historical frames. In this context, FFNet [153] was designed to leverage temporal coherence by extracting the first-order derivative of feature flow from previous infrastructure frames to predict future features and mitigate asynchrony. Although this first-order expansion method was effective over very short time intervals, it was found to have inherent limitations. In SyncNet [154], a dual-branch pyramid LSTM network was employed to infer real-time features and cooperative attention. Similarly, the HFP module [156] was developed to generate predictions of the current frame through a series of multidimensional convolutional mixing operations. However, these methods have not effectively addressed the challenges introduced by irregular time delays. As an alternative, each collaboration message was treated as an irregular sample in CoBEVFlow [155], and BEV flow map was estimated using historical frames for asynchronous feature alignment, although reliance on past data for latency compensation remained. In contrast, this dependency was bypassed in V2X-PC [50] by directly utilizing low-level coordinate information in point clusters to predict their positions at the current timestamp.
Re-encoding received features with delay time has been explored as an effective approach for mitigating communication delays. For instance, in V2X-ViT [59], time delay was addressed through a Delay-aware Positional Encoding module, in which sinusoidal functions conditioned on delay time and channel information were used to initialize a learnable embedding. Similarly, in V2VNet [54], a CNN was employed to manage latency by incorporating the received intermediate feature, relative 6DoF pose, and delay time to produce a time-delay-compensated representation. These methods assumed that latency was known. However, latency was typically computed using timestamps from different agents’ clocks, which could introduce deviations and result in inaccurate measurements. To address this issue, a multi-anchor-based subgraph searching algorithm was introduced in FreeAlign [107] to identify a common subgraph between salient-object graphs across different agents, enabling estimation of clock deviations and determination of the precise time difference between collaborative messages. This approach ensured accurate temporal alignment without relying on external clock synchronization signals.
@@ V-D1 Data Security
While cooperative multi-agent systems offer significant promise, they also present security risks due to the potential for malicious or unreliable communication between agents [180]. Malicious agents may tamper with shared packets or generate forged data. The ego vehicle is unable to independently verify missed events or confirm the authenticity of received data [181]. Imperceptible perturbations introduced by adversaries can substantially alter perception outputs, compromising system integrity. Although adversarial training has been used to address these threats, it adds training overhead and often fails to generalize to novel attacks [157]. As a solution, ROBOSAC [158] was proposed, allowing ego vehicles to intelligently select trustworthy collaborators. By detecting divergences in adversarial messages and aligning with reliable collaborators, ROBOSAC generalizes to unseen adversarial strategies and offers a flexible trade-off between performance and computational cost.
@@
To ensure data security, assessing the quality of received information and detecting erroneous data before processing is crucial. Probabilistic modelling techniques have been utilized to estimate the trustworthiness of connected agents based on data consistency [159]. This approach applied a Bayes filter to quantify consistency, allowing the detection of forged information by analyzing the correlation between data consistency and agent reliability. However, this method was limited in detecting malicious agents, as it relied on substantial redundancy, typically requiring at least three conflicting sources, to identify attackers effectively.
@@
FDII [160] was proposed to address spoofing attacks by leveraging LiDAR data from neighbouring vehicles to detect tampered points in a victim’s scan. Discrepancies were identified by comparing altered and unaltered scans, and attacks were classified using a decision tree algorithm, which was also used to update the unsafe areas associated with each vehicle. This approach was shown to be effective in identifying attack types and reconstructing vulnerable regions. However, decision trees were observed to struggle with complex data relationships and often exhibited bias toward the majority class in imbalanced datasets, potentially resulting in suboptimal classification of minority classes.
@@
To counter data fabrication attacks, the Collaborative Anomaly Detection (CAD) model [161] was proposed, enabling benign vehicles to collectively detect malicious fabrications by sharing and merging occupancy maps. These maps, represented as fine-grained polygons, were shown to offer greater precision and flexibility compared to grid-based alternatives. Motion estimation was incorporated to track objects across frames, with motion data attached to the maps. Consistency checks were performed to identify discrepancies among synchronized maps and to compare individual perception results with the merged map, thereby raising alerts for conflicted regions indicative of potential anomalies. Although CAD was demonstrated to be effective in real-world scenarios, its efficacy was found to depend on the presence of at least one benign CAV monitoring the attacked region.
@@
Malicious Agent Detection (MADE) [162] was developed to identify malicious agents using two key statistical metrics: match loss and collaborative reconstruction loss. Match loss was used to quantify discrepancies in bounding box proposals between the ego agent and the inspected agent, while collaborative reconstruction loss was employed to measure the consistency of feature maps when the two agents collaborated. These metrics were applied either independently or jointly to construct a multi-test with a controlled false detection rate. Although this approach provided valuable insights, its reliance on conformal p-values computed from calibration sets was based on the assumption that the ego agent was benign—an assumption that may not always hold in real-world scenarios.